OpenStudy (anonymous):

Find the zeros of f(x)=x^2-2x-8

OpenStudy (anonymous):

There is a website called mathway.com that you can usually type problems like this into and it'll automatically compute the answer for you.

OpenStudy (anonymous):

Factor the expression gives \[x^2-2x-8 = (x+2)(x-4)\] And we notice that for this expression to give 0, either (x+2) has be to 0, or (x-4) has to be 0. Hence the zeros are -2 or 4

OpenStudy (anonymous):

Okay thank you. So you telling me that the zeros of this functions is -2 or 4?

OpenStudy (anonymous):

There are 2 zeros, and they are -2 AND 4 :)

OpenStudy (anonymous):

How do I figure it how? how do you work this problem?

OpenStudy (anonymous):

okay thank you . but how did you get -2 and 4

OpenStudy (anonymous):

One way is to write: (x )(x ) Then you think about which 2 numbers multiply together gives -8 and add up together gives -2

OpenStudy (anonymous):

So you might think 1 and -8 multiply each other gives -8, but add up doesn't give -2, then you take another couple of numbers and check if it works, then you find out the only couple that works are 2 and -4

OpenStudy (anonymous):

Okay thank you. what if I had a simpler problem like find the zero(s) of f(x)= 3x-2.. How would you work this problem??

OpenStudy (anonymous):

You want 3x -2 = 0 in this case. Add 2 to both sides to get 3x = 2 and then divide both sides by 3 to get your answer.

OpenStudy (anonymous):

Which would be 2/3??

OpenStudy (anonymous):

@brooke_tessa Basically, finding zeroes is just another way of saying: when x=?, the whole expression equals to 0?

OpenStudy (anonymous):

it would just be like the following: f(x)=3x-2 *set it to 0 0=3x-2 x=2/3

OpenStudy (anonymous):

Yes...nice job!

OpenStudy (anonymous):

OKAY! Thank y'all. so if I had a problem like f(x)= -4x + 6 -6= -4x = 6/4 = 3/2 ???? IS THAT RIGHT??

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

OKAY!!! THANK YALL SO MUCH FOR THE HELP! I UNDERSTAND (: